OpenStudy (anonymous):

complete the square z^2=16z-64

OpenStudy (anonymous):

try putting everything on one side. \[z^2 -16z + 64 = 0\]

OpenStudy (anonymous):

can you see it now?

OpenStudy (anonymous):

i got that part

OpenStudy (anonymous):

You need to add half of the coefficient of z and square it (do it to both sides).\[z^2-16z=-64 \rightarrow z^2-16z+(-\frac{16}{2})^2=-64+(-\frac{16}{2})^2\]

OpenStudy (anonymous):

\[z^2-16z+(-8)^2=-64+(-8)^2\]You can read off the left now as:\[(z-8)^2\]and the right is simply 0, so your equation becomes,\[(z-8)^2=0\]

OpenStudy (anonymous):

Which means one solution for z, z=8.

OpenStudy (anonymous):

its suppose to be two answers so would it be 8,-8??

OpenStudy (anonymous):

Well, not in this case, since the right-hand side ended up being zero. When you take the square root of both sides, z-8 = +/- 0 = 0 so z = 8. If you'd had, (z-8)^2 = 7, for example, THEN you'd have two solutions:\[(z-8)^2=7 \rightarrow z-8=\pm \sqrt{7} \rightarrow z=8 \pm \sqrt{7}\]

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

welcome :) hope you pass!